OpenStudy (anonymous):

The altitude of a triangle is increasing at a rate of 2000 centimeters/minute while the area of the triangle is increasing at a rate of 2000 square centimeters/minute. At what rate is the base of the triangle changing when the altitude is 8000 centimeters and the area is 96000 square centimeters?

OpenStudy (michele_laino):

from the text of your problem, we can write: \[\frac{ dh }{ dt }=2000, \frac{ dS }{ dt }=2000\] where h is the altitude of your triangle, and S is the area of that triangle, t is time Now, we have 2S=b*h, where b is the base of your triangle. Differentiating the preceeding formula for S, we get: \[2\frac{ dS }{ dt }=\frac{ db }{ dt }*h+b*\frac{ dh }{ dt }\] Please try to substitute your numerical data into the above equation, and you will get the rate for b. Please write youe answer

OpenStudy (michele_laino):

I got db/dt=570 cm/min